# Post-measurement Quantum Monte Carlo

### v1

@@ Abstract
We show how the effects of large numbers of measurements on many-body quantum ground and thermal states can be studied using Quantum Monte Carlo (QMC). Density matrices generated by measurement in this setting feature products of many local nonunitary operators, and by expanding these density matrices as sums over operator strings we arrive at a generalized stochastic series expansion (SSE). Our ‘post-measurement SSE’ is based on importance sampling of operator strings contributing to a measured thermal density matrix. We demonstrate our algorithm by probing the effects of measurements on the spin-$1/2$ Heisenberg antiferromagnet on the square lattice. Thermal states of this system have $\mathrm{SU}(2)$ symmetry, and we preserve this symmetry by measuring $\mathrm{SU}(2)$ symmetric observables. We identify classes of post-measurement states for which correlations can be calculated efficiently, as well as states for which $\mathrm{SU}(2)$ symmetric measurements generate a QMC sign problem when working in any site-local basis. For the first class, we show how deterministic loop updates can be leveraged. Using our algorithm we demonstrate the creation of long-range Bell pairs and symmetry-protected topological order, as well as the measurement-induced enhancement of antiferromagnetic correlations. The method developed in this work opens the door to scalable experimental probes of measurement-induced collective phenomena, which require numerical estimates for the effects of measurements.
@@ I Introduction
The measurements we consider preserve the $\mathrm{SU}(2)$ symmetry of the state. By measuring the total spin on nearby pairs of spin-$1/2$ degrees of freedom we generate either local singlet or triplet degrees of freedom. While a singlet outcome necessarily disentangles the measured spins from the rest of the system, a triplet outcome generates an effective spin-$1$ degree of freedom, and these local three-level systems can remain entangled with their surroundings.
@@
We choose this setup to illustrate a variety of features of the post-measurement SSE. For generic measurement outcomes, and when working with a basis of product states, the post-measurement SSE features a sign problem; the sign problem which we generate has a similar origin those familiar in the study of frustrated antiferromagnets [27]. The presence of a sign problem in the post-measurement SSE is to be contrasted with its absence in the standard (thermal) SSE for our model. For post-measurement states whose SSEs do not suffer from a sign problem, we show that the operator strings contributing to the density matrix can be sampled using loop updates, thereby allowing for efficient calculations of observable expectation values. We use these states (which, in experiment, would arise from post-selection) to explore the implications of the $\mathrm{SU}(2)$ symmetry for post-measurement correlations.
@@
First, we study the generation of long-range singlet states and symmetry-protected topological (SPT) [48] states using local measurements. We show also that the symmetry of the system allows for a straightforward calculation of mixed-state entanglement measures using QMC. Following this, we show that the interplay between correlations in the initial thermal density matrix and (non-overlapping) local measurements can enhance long-range correlations. Using these measurements we probe an $\mathrm{O}(3)$ quantum critical point, and provide evidence that in certain post-selected states the correlations decay parametrically slower than in the thermal state. This behavior is expected for a quantum state whose correlations realize the extraordinary-log boundary universality class of the $\mathrm{O}(3)$ model [49, 17].
@@
This work is organized as follows. In Sec. II we provide background on the analysis of measurement-induced phenomena, the model studied in this work, and the standard SSE. The post-measurement SSE is introduced in Sec. III; there we discuss how the effects of both projective and weak measurements can be captured within a QMC calculation. In Sec IV we demonstrate the creation of long-range singlet and SPT states. In Sec. V we study the measurement-induced enhancement of antiferromagnetic correlations in post-selected states. We discuss our results as well as open questions in Sec. VI.
@@ II.1 Post-measurement states
The effects of measurements on ground and thermal states can be understood via the quantum-classical mapping [26, 16]. On writing the thermal density matrix as a Euclidean spacetime path integral, spatial (equal-time) correlations in this state become correlations within a $d$-dimensional surface of the $(d+1)$-dimensional spacetime. We refer to this surface as ‘$\tau=0$’, where the imaginary time $\tau\in[0,\beta]$ runs around a thermal circle of circumference $\beta$. When calculating post-measurement correlation functions using this formalism, measurements appear as defects in the $\tau=0$ surface. This follows simply from the fact that in the post-measurement density matrix $\rho^{\bm{s}}\sim P^{\bm{s}}e^{-\beta H}P^{\bm{s}}$ the projectors $P^{\bm{s}}$ appear on either side of the imaginary-time evolution operator $e^{-\beta H}$, rather than ‘in the middle of’ this evolution.
@@ II.2 Model
where $j=0,\ldots,(N-1)$ labels sites and NN is the set of nearest neighbors. The operators $\bm{S}_{j}=(S^{x}_{j},S^{y}_{j},S^{z}_{j})$ satisfy the standard commuation relations $[S^{a}_{j},S^{b}_{k}]=i\delta_{jk}\epsilon^{abc}S^{c}_{j}$, and the initial unmeasured density matrix is $\rho=e^{-\beta H}/\text{Tr}e^{-\beta H}$. We use periodic boundary conditions along both directions, with an even number of lattice sites along each to prevent frustration. In systems having length $N_{x}$ in the $x$ direction and $N_{y}$ in the y direction, our convention is that the coordinates of site $j$ are $(x_{j},y_{j})$ with $x_{j}=j\bmod{N_{y}}$ and $y_{j}=j-x_{j}N_{y}$. The couplings $J_{jk}\geq 0$ are unfrustrated, so the ground state of $H$ is a singlet under global $\mathrm{SU}(2)$ transformations [50].
[…]
@@
For all $\beta$ the thermal density matrix has a weak symmetry $U\rho U^{-1}=\rho$ under global $\mathrm{SU}(2)$ rotations $U$, and because of the $\mathrm{SU}(2)$ symmetry of $\bm{S}_{j}\cdot\bm{S}_{k}$ this feature is shared with the post-measurement states. Similarly, for $\beta\to\infty$ both the unmeasured and post-measurement ground states have strong $\mathrm{SU}(2)$ symmetry, $\rho=U\rho=\rho U$, since they are total singlets.
@@ II.3 Stochastic Series Expansion
The SSE is based on (i) an expansion of $e^{-\beta H}$ as a sum over operator strings, where each string is a product of operators contributing to $H$, and (ii) an importance sampling procedure of these operator strings. As our computational basis we choose eigenstates |$\ket{\alpha}$ of all $S^{z}_{j}$ operators, writing the partition function as
$Z=\text{Tr}[e^{-\beta H}]=\sum_{\alpha}\sum_{n=0}^{\infty}\frac{(-\beta)^{n}}{n!}\braket{\alpha}{H^{n}}{\alpha}.$ (10)
Next, we decompose $H$ into a sum over terms that do not cause ‘branching’ in the computational basis. This means that the action of each term in the expansion on a state |$\ket{\alpha}$ generates a unique new basis state |$\ket{\alpha^{\prime}}$, rather than a linear combination of such states. A standard choice is to decompose the two site Heisenberg operator $\bm{S}_{j}\cdot\bm{S}_{k}$ into its diagonal and off-diagonal parts [27]. We define
@@
Having defined our operator strings we can assign statistical weights. The weight corresponding to basis state |$\ket{\alpha}$ and operator string $\mathcal{H}_{\bm{d}}$ is $\mathcal{C}^{H}_{\bm{d}}\braket{\alpha}{\mathcal{H}}{\alpha}$ where
@@
and the product of couplings $J_{(jk)_{p}}$ only runs over $d=1$ and $d=2$ type operators. The specific prefactors of $O^{H}_{d,(jk)}$ have been chosen so that the above matrix element is a boolean variable: $\braket{\alpha}{\mathcal{H}}{\alpha}=1$ for allowed configurations and $\braket{\alpha}{\mathcal{H}}{\alpha}=0$ otherwise. Note that on a bipartite lattice $\braket{\alpha}{\mathcal{H}}{\alpha}\neq 0$ only when $L_{2}$, the number of ‘hops’, is even. Therefore, all contributing weights are non-negative. The combinatorial prefactor in $\mathcal{C}^{H}_{\bm{d}}$ is the ratio of (i) $1/(L_{1}+L_{2})!$ from the Taylor expansion and (ii) (${L\choose L_{1}+L_{2}}$, where (ii) compensates for the (${L\choose L_{1}+L_{2}}$ ways of arranging non-identity operators in the length $L$ operator string. The full partition function in the fixed length scheme is
$Z=\sum_{\alpha\bm{d}}\mathcal{C}^{H}_{\bm{d}}\braket{\alpha}{\mathcal{H}_{\bm{d}}}{\alpha}.$ (16)
In order to sample basis states |$\ket{\alpha}$ and operator strings $\mathcal{H}_{\bm{d}}$ according to the weights $\mathcal{C}^{H}_{\bm{d}}\braket{\alpha}{\mathcal{H}}{\alpha}$ we use two kinds of update, (i) ‘diagonal’ moves, in which $d=0$ operators are replaced by $d=1$ operators, and vice versa, and (ii) ‘off-diagonal’ moves in which $d=1$ and $d=2$ operators replace one another. For the latter we can use operator loop updates [52].
The different contributions to $Z$ can be visualized in terms of progagated states |$\ket{\alpha(p)}$ with $p=0,\ldots,L-1$ defined modulo $L$. Each operator string $\mathcal{H}_{\bm{d}}$, here of the form in Eq. 14, can be factorized as $\mathcal{H}_{\bm{d}}=\mathcal{H}^{\geq p}_{\bm{d}}\mathcal{H}^{<p}_{\bm{d}}$, where $\mathcal{H}^{<p}_{\bm{d}}$ consists of the first $p$ operators in the string, and $\mathcal{H}^{\geq p}_{\bm{d}}$ consists of the final $L-p$. Here the propagated states have the form |$\ket{\alpha(p)}=\mathcal{H}^{<p}_{\bm{d}}\ket{\alpha}$, with |$\ket{\alpha}=\ket{\alpha(0)}=\ket{\alpha(L)}$ for $\braket{\alpha}{\mathcal{H}_{\bm{d}}}{\alpha}=1$. These expressions lead to an interpretation of the discrete index $p$ as a label for successive ‘slices’ of the (continuous) imaginary time in the Euclidean spacetime path integral.
Before moving on to the post-measurement SSE, we discuss the problem of generating measurement outcomes according to the Born rule, i.e. ‘sampling’ from a quantum state via measurement. Consider an experiment where $S^{z}_{j}$ is measured (projectively) on a set of sites $B$; the probability that our measurements return the outcomes |$\ket{\alpha_{B}}$ is
@@
where |$\ket{\alpha}=\ket{\alpha_{A},\alpha_{B}}$ is a basis state of the full system, with $A$ the complement of $B$. From this expression it is clear that the SSE automatically generates configurations of ‘measurement outcomes’ |$\ket{\alpha_{B}}$ according to the Born rule. This example is particularly simple because the measurement basis matches the computational basis. We now turn to the more general setting where these two bases are different, and we will also show how to calculate expectation values in the states generated by measurement.
@@ III.1 Projective measurements
where the superscript indicates the outcome for the total spin $s_{(jk)}$ on the measured bond (00 for singlet and $1$ for triplet). In the singlet case the operators are
@@
$Z^{\bm{s}}=\sum_{\alpha\bm{a}\bm{b}\bm{c}}\mathcal{C}^{H}_{\bm{a}}\mathcal{C}^{\bm{s}}_{\bm{b}}\mathcal{C}^{\bm{s}}_{\bm{c}}\braket{\alpha}{\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}}{\alpha}.$ (25)
The post-measurement SSE is based on Eq. (25). Note that the operator string $\mathcal{P}_{\bm{b}}^{\bm{s}}\mathcal{H}_{\bm{a}}\mathcal{P}_{\bm{c}}^{\bm{s}}$, which has length $L+2|B|$, does not cause branching of basis states. From our definitions of $O^{s_{(jk)}}_{d,(jk)}$, the matrix elements $\braket{\alpha}{\mathcal{P}_{\bm{b}}^{\bm{s}}\mathcal{H}_{\bm{a}}\mathcal{P}_{\bm{c}}^{\bm{s}}}{\alpha}=0,1$ remain boolean variables, and the coefficient $\mathcal{C}^{H}_{\bm{a}}$ has the same form as in the standard SSE discussed in the previous section. Propagated states |$\ket{\alpha(p)}$ can here be defined, in analogy with those appearing in the thermal SSE, as the basis states generated by acting with successive operators in $\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}$ on |$\ket{\alpha}$, and we illustrate an example configuration in Fig. 3.
@@
In Fig. 3 we illustrate an operator string contributing to the post-measurement SSE. Expectation values must be evaluated at imaginary time ‘$\tau=0$’, i.e. between the two measurement strings as indicated above. This situation should be contrasted with the thermal SSE, where imaginary-time translation symmetry of operator strings means that expectation values can be evaluated as averages over ‘time slices’ $\tau$. This difference arises from the fact that measurements explicitly break the time-translation symmetry of the imaginary-time path integral.
Figure 3: Example post-measurement SSE configuration for a spin chain, with all propagated states shown. Dashed lines indicate that the propagated state on a site is unchanged with respect to the previous step. Open (filled) circles denote $\ket{\downarrow}$ ($\ket{\uparrow}$) states, while open (solid) bars represent diagonal $O_{1,(jk)}$ (off-diagonal $O_{2,(jk)}$) operators, and the absence of a bar represents an identity operator. The configuration displayed consists of $L_{1}=6$ diagonal operators located after $p=0,1,7,11$ and $L_{2}=6$ off-diagonal operators after $p=0,2,4,6,8,11$. The post-measurement correlations of diagonal observables are evaluated, with respect to |$\ket{\alpha(p=0)}=\ket{\alpha}$, in the $p=0$ imaginary time slice.
@@ III.2 Weak measurements
From this step on, the analysis is essentially identical to projective measurements. Inserting the decompositions of operator strings we find precisely that the partition function $Z^{\bm{s}}_{\mu}$ differs from $Z^{\bm{s}}$ ($=Z^{\bm{s}}_{\infty}$) in Eq. (25) only by the replacement of coefficients $\mathcal{C}^{\bm{s}}_{\bm{d}}$ ($=\mathcal{C}^{\bm{s}}_{\bm{d},\infty}$) with $\mathcal{C}^{\bm{s}}_{\bm{d}}$. The values of these coefficients can be determined from Eq. (31); for example, in the cases where $\bm{s}=\bm{0}$ and $\bm{s}=\bm{1}$ on all measured pairs of sites, these coefficients are respectively equal to
@@
For consistency, we have here described the weak measurement scheme along similar lines to the projective measurement scheme. Note, however, that we can also write $W^{\bm{s}}\propto e^{-\mu M^{\bm{s}}/2}$, where $M^{\bm{s}}$ is a sum of local terms. In this notation the partition function involves a product of $e^{-\mu M^{\bm{s}}}$ and $e^{-\beta H}$. Just as $e^{-\beta H}$ is expanded in powers of $\beta$, the operator $e^{-\mu M^{\bm{s}}}$ can be expanded in powers of $\mu$. This leads to an algorithm where we view $M^{\bm{s}}$ as the generator of ‘imaginary time’ evolution for an interval $\mu$. We detail this algorithm in Appendix A, and use this in practice for our finite $\mu$ numerical calculations. Note that this algorithm allows for a general description of correlations in systems described by a combination of different imaginary time evolutions. For example, it also allows for the investigation of strange correlators [53], which were previous calculated using QMC in Ref [54].
Figure 4: Allowed vertices in the measurement string for (upper row) $s=0$ outcomes and (lower row) $s=1$ outcomes. Open and filled circles represent $\ket{\downarrow}$ and $\ket{\uparrow}$ states, respectively, while open and filled boxes represent (non-identity) diagonal $O^{s}_{1,(jk)}$ and off-diagonal $O^{s}_{2,(jk)}$ operators. Diagonal operators act on anti-parallel spins for $s=0$, whereas they act on parallel spins for $s=1$.

### v2

@@ Abstract
We show how the effects of large numbers of measurements on many-body quantum ground and thermal states can be studied using Quantum Monte Carlo (QMC). Density matrices generated by measurement in this setting feature products of many local nonunitary operators, and by expanding these density matrices as sums over operator strings we arrive at a generalized stochastic series expansion (SSE). Our ‘post-measurement SSE’ is based on importance sampling of operator strings contributing to a measured thermal density matrix. We demonstrate our algorithm by probing the effects of measurements on the spin-$1/2$ Heisenberg antiferromagnet on the square lattice. Thermal states of this system have $\mathrm{SU}(2)$ symmetry, and at first we preserve this symmetry by measuring $\mathrm{SU}(2)$ symmetric observables. We identify classes of post-measurement states for which correlations can be calculated efficiently, as well as states for which $\mathrm{SU}(2)$ symmetric measurements generate a QMC sign problem when working in any site-local basis. For the first class, we show how deterministic loop updates can be leveraged. Using our algorithm we demonstrate the creation of long-range Bell pairs and symmetry-protected topological order, as well as the measurement-induced enhancement of antiferromagnetic correlations. We then study the effects of measuring the system in a basis where the standard (unmeasured) SSE is sign-free: For measurement schemes with this property, we can calculate correlations in all post-measurement states without a sign problem. The method developed in this work opens the door to scalable experimental probes of measurement-induced collective phenomena, which require numerical estimates for the effects of measurements.
@@ I Introduction
Our numerical method can be applied to study the effects of measurements on ground and thermal states provided there exists a basis for which the standard (unmeasured) SSE is free of a QMC sign problem. Measurements of operators that are diagonal in that sign-free basis can then be introduced to construct post-measurement SSEs that are themselves sign-free for all possible measurement outcomes. However, measurements of operators that are not diagonal in the sign-free basis generically lead to post-measurement SSEs that do have sign problems.
@@
Through the majority of this work, the measurements we study preserve the $\mathrm{SU}(2)$ symmetry of the state. By measuring the total spin on nearby pairs of spin-$1/2$ degrees of freedom we generate either local singlet or triplet degrees of freedom. While a singlet outcome necessarily disentangles the measured spins from the rest of the system, a triplet outcome generates an effective spin-$1$ degree of freedom, and these local three-level systems can remain entangled with their surroundings.
@@
We choose this setup to illustrate a variety of features of the post-measurement SSE. For generic measurement outcomes, and when working with a basis of product states, the post-measurement SSE features a sign problem; the sign problem which we generate has a similar origin those familiar in the study of frustrated antiferromagnets [27]. The presence of a sign problem in the post-measurement SSE is to be contrasted with its absence in the standard (thermal) SSE for our model. For the subset of post-measurement states whose SSEs do not suffer from a sign problem, we show that the operator strings contributing to the density matrix can be sampled using loop updates, thereby allowing for efficient calculations of observable expectation values. We use these states (which, in experiment, would arise from post-selection) to explore the implications of the $\mathrm{SU}(2)$ symmetry for post-measurement correlations.
@@
First, we study the generation of long-range singlet states and symmetry-protected topological (SPT) [48] states using local measurements. We show also that the symmetry of the system allows for a straightforward calculation of mixed-state entanglement measures using QMC. Following this, we show that the interplay between correlations in the initial thermal density matrix and (non-overlapping) local measurements can enhance long-range correlations. Using these measurements we probe an $\mathrm{O}(3)$ quantum critical point, and provide evidence that in certain post-selected states the correlations decay parametrically slower than in the ground state. This behavior is expected for a quantum state whose correlations realize the extraordinary-log boundary universality class of the $\mathrm{O}(3)$ model [49, 17].
Finally, we study the measurement of standard Pauli operators. In this case, when all measurements are in the same basis, for all possible measurement outcomes the post-measurement SSE has no sign problem. Because of this we can efficiently calculate averaged properties of the ensemble of post-measurement states. At the $\mathrm{O}(3)$ quantum critical point, such measurements are expected to behave as a relevant perturbation [16]; we verify numerically that the critical correlations are destroyed when we measure the spins on one of the two sublattices.
@@
This work is organized as follows. In Sec. II we provide background on the analysis of measurement-induced phenomena, the model studied in this work, and the standard SSE. The post-measurement SSE is introduced in Sec. III; there we discuss how the effects of both projective and weak measurements can be captured within a QMC calculation. In Sec IV we demonstrate the creation of long-range singlet and SPT states. In Sec. V we study the measurement-induced enhancement of antiferromagnetic correlations in post-selected states. Then, in Sec. VI, we study the full ensemble of post-measurement states generated by measurements of spin operators. We discuss our results as well as open questions in Sec. VII.
@@ II.1 Post-measurement states
The effects of measurements on ground and thermal states can be understood via the quantum-classical mapping [26, 16]. On writing the thermal density matrix as a Euclidean spacetime path integral, spatial (equal-time) correlations in this state become correlations within a $d$-dimensional surface of the $(d+1)$-dimensional spacetime. We refer to this surface as $\tau=0$’, where the imaginary time $\tau\in[0,\beta]$ runs around a thermal circle of circumference $\beta$. When calculating post-measurement correlation functions using this formalism, measurements appear as defects in the $\tau=0$ surface. This follows simply from the fact that in the post-measurement density matrix $\rho^{\bm{s}}\sim P^{\bm{s}}e^{-\beta H}P^{\bm{s}}$ the projectors $P^{\bm{s}}$ appear on either side of the imaginary-time evolution operator $e^{-\beta H}$, rather than ‘in the middle of’ this evolution.
@@ II.2 Model
where $j=0,\ldots,(N-1)$ labels sites and NN is the set of nearest neighbors. The operators $\bm{S}_{j}=(S^{x}_{j},S^{y}_{j},S^{z}_{j})$ satisfy the standard commuation relations $S^{a}_{j},S^{b}_{k}]=i\delta_{jk}\epsilon^{abc}S^{c}_{j}$, and the initial unmeasured density matrix is $\rho=e^{-\beta H}/\text{Tr}e^{-\beta H}$. We use periodic boundary conditions along both directions, with an even number of lattice sites along each to prevent frustration. In systems having length $N_{x}$ in the $x$ direction and $N_{y}$ in the y direction, our convention is that the coordinates of site $j$ are $(x_{j},y_{j})$ with $x_{j}=j\bmod{N_{y}}$ and $y_{j}=j-x_{j}N_{y}$. The couplings $J_{jk}\geq 0$ are unfrustrated, so the ground state of $H$ is a singlet under global $\mathrm{SU}(2)$ transformations [50].
$J_{2}$$J_{1}$
@@
For all $\beta$ the thermal density matrix has a weak symmetry $U\rho U^{-1}=\rho$ under global $\mathrm{SU}(2)$ rotations $U$, and because of the $\mathrm{SU}(2)$ symmetry of $\bm{S}_{j}\cdot\bm{S}_{k}$ this feature is shared with the post-measurement states. Similarly, for $\beta\to\infty$ both the unmeasured and post-measurement ground states have strong $\mathrm{SU}(2)$ symmetry, $\rho=U\rho=\rho U$, since they are total singlets. Note that, in Sec. VI, we will also study the effects of symmetry-breaking measurements.
@@ II.3 Stochastic Series Expansion
[…]
$Z=\text{Tr}[e^{-\beta H}]=\sum_{\alpha}\sum_{n=0}^{\infty}\frac{(-\beta)^{n}}{n!}\braket{\alpha|H^{n}|\alpha}.$ (10)
[…]
@@
[…]
@@
and the product of couplings $J_{(jk)_{p}}$ only runs over $d=1$ and $d=2$ type operators. The specific prefactors of $O^{H}_{d,(jk)}$ have been chosen so that the above matrix element is a boolean variable: $\braket{\alpha|\mathcal{H}|\alpha}=1$ for allowed configurations and $\braket{\alpha|\mathcal{H}|\alpha}=0$ otherwise. Note that on a bipartite lattice $\braket{\alpha|\mathcal{H}|\alpha}\neq 0$ only when $L_{2}$, the number of ‘hops’, is even. Therefore, all contributing weights are non-negative. The combinatorial prefactor in $\mathcal{C}^{H}_{\bm{d}}$ is the ratio of (i) $1/(L_{1}+L_{2})!$ from the Taylor expansion and (ii) $L\choose L_{1}+L_{2}$}, where (ii) compensates for the $L\choose L_{1}+L_{2}$} ways of arranging non-identity operators in the length $L$ operator string. The full partition function in the fixed length scheme is
$Z=\sum_{\alpha\bm{d}}\mathcal{C}^{H}_{\bm{d}}\braket{\alpha|\mathcal{H}_{\bm{d}}|\alpha}.$ (16)
@@
[…]
$\braket{\alpha_{B}|\text{Tr}_{A}[\rho]|\alpha_{B}}=Z^{-1}\sum_{\alpha_{A}\bm{d}}\mathcal{C}^{H}_{\bm{d}}\braket{\alpha|\mathcal{H}|\alpha}$ (17)
[…]
@@ III.1 Projective measurements
where the superscript indicates the outcome for the total spin $s_{(jk)}$ on the measured bond (0 for singlet and $1$ for triplet). In the singlet case the operators are
@@
$Z^{\bm{s}}=\sum_{\alpha\bm{a}\bm{b}\bm{c}}\mathcal{C}^{H}_{\bm{a}}\mathcal{C}^{\bm{s}}_{\bm{b}}\mathcal{C}^{\bm{s}}_{\bm{c}}\braket{\alpha|\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}|\alpha}.$ (25)
[…]
@@
$\braket{A}^{\bm{s}}=\frac{1}{Z^{\bm{s}}}\sum_{\alpha\bm{a}\bm{b}\bm{c}}\mathcal{C}^{H}_{\bm{a}}\mathcal{C}^{\bm{s}}_{\bm{b}}\mathcal{C}^{\bm{s}}_{\bm{c}}\braket{\alpha|\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}|\alpha}\braket{\alpha|A|\alpha}.$ (26)
In Fig. 3 we illustrate an operator string contributing to the post-measurement SSE. Expectation values must be evaluated at imaginary time $\tau=0$’, i.e. between the two measurement strings as indicated above. This situation should be contrasted with the thermal SSE, where imaginary-time translation symmetry of operator strings means that expectation values can be evaluated as averages over ‘time slices’ $\tau$. This difference arises from the fact that measurements explicitly break the time-translation symmetry of the imaginary-time path integral.
@@ III.2 Weak measurements
From this step on, the analysis is essentially identical to projective measurements. Inserting the decompositions of operator strings we find precisely that the partition function $Z^{\bm{s}}_{\mu}$ differs from $Z^{\bm{s}}$ ($Z^{\bm{s}}_{\infty}$) in Eq. (25) only by the replacement of coefficients $\mathcal{C}^{\bm{s}}_{\bm{d}}$ ($\mathcal{C}^{\bm{s}}_{\bm{d},\infty}$) with $\mathcal{C}^{\bm{s}}_{\bm{d}}$. The values of these coefficients can be determined from Eq. (31); for example, in the cases where $\bm{s}=\bm{0}$ and $\bm{s}=\bm{1}$ on all measured pairs of sites, these coefficients are respectively equal to
@@
For consistency, we have here described the weak measurement scheme along similar lines to the projective measurement scheme. Note, however, that we can also write $W^{\bm{s}}\propto e^{-\mu M^{\bm{s}}/2}$, where $M^{\bm{s}}$ is a sum of local terms. In this notation the partition function involves a product of $e^{-\mu M^{\bm{s}}}$ and $e^{-\beta H}$. Just as $e^{-\beta H}$ is expanded in powers of $\beta$, the operator $e^{-\mu M^{\bm{s}}}$ can be expanded in powers of $\mu$. This leads to an algorithm where we view $M^{\bm{s}}$ as the generator of ‘imaginary time’ evolution for an interval $\mu$. We detail this algorithm in Appendix B, and use this in practice for our finite $\mu$ numerical calculations. Note that this algorithm allows for a general description of correlations in systems described by a combination of different imaginary time evolutions. For example, it also allows for the investigation of strange correlators [53], which were previous calculated using QMC in Ref [54].
[…]